Parth (parthkohli):

\[79^{79} \equiv N \pmod{100}\]\(N\) is a two digit number.

Parth (parthkohli):

Fermat's Little Theorem?

Parth (parthkohli):

\[79 \equiv 79 \equiv 0 \pmod{79} \]Which is kinda obvious.

terenzreignz (terenzreignz):

And so we meet again.

Parth (parthkohli):

Euler's Theorem says that \(\phi(79) = 79\) T_T

Parth (parthkohli):

Yeah ;-)

terenzreignz (terenzreignz):

Yeah? I was talking to the modular maths problem XD LOL Jk

terenzreignz (terenzreignz):

So anyway, 79^5 = -1(mod 100)

Parth (parthkohli):

How do you know again...

terenzreignz (terenzreignz):

79^2 = 6241 = 41(mod 100) 79^3 = (41)(79)(mod 100) = 39(mod 100) 79^4 = (39)(79)(mod 100) = 81(mod 100) 79^5 = (81)(79)(mod 100) = 99(mod 100) = -1(mod 100)

Parth (parthkohli):

So \(79^{10} \equiv 1 \pmod{100}\)?

terenzreignz (terenzreignz):

Yeah, that's true....

terenzreignz (terenzreignz):

Meaning 79^70 = 1(mod 100), too :D

Parth (parthkohli):

Okay, oh.

Parth (parthkohli):

So does that mean \(79^{80} \equiv 1 \pmod {100}\)

Parth (parthkohli):

Ah, yeah.

terenzreignz (terenzreignz):

Yes... But it's 79^79 you want, right? 79^75 = -1(mod 100)

Parth (parthkohli):

\[79^{70} \times 79^{5} \times 79^{4} \pmod{100}\]AH!

terenzreignz (terenzreignz):

mhmm... 79^70 is 1 79^5 is -1 79^4, refer to the process we did earlier...

Parth (parthkohli):

\[1 \times -1 \times 81 \pmod{100}\]

terenzreignz (terenzreignz):

Yeah, that seems right :D

Parth (parthkohli):

\[19\pmod{100}\]

terenzreignz (terenzreignz):

Bingo... :|

Parth (parthkohli):

:-D